OpenStudy (anonymous):

If √6 is the geometric mean between 4 and another number, then the number is 1.5 9 24

OpenStudy (anonymous):

1.5

OpenStudy (anonymous):

√(4*x) = √6 where x represents the unknown number. Now, solve for the unknown number x but taking the square of each side of the equal sign to get rid of the square root sign: [√(4*x)]^2 = [√6]^2 4*x = 6 Solve for x: x = 6/4 x = 3/2 3/2 = 1.5